# Hopfield-Fenchel-Young Networks: A Unified Framework for Associative Memory Retrieval

### v2

@@ 6.2 Sequential recall
Algorithm 3 Sequential recall using SparseMAP with sequential $2$-subsets. $t>0$ denotes the transition score and $\omega\geq 1$ is a coefficient which promotes sequential order by boosting the emission score of the last recalled pattern.
1:$\bm{X}$, $\bm{q}$, $\lambda$, $\tau$, $\omega$, $t$, $\beta$, $T$
2:$\bm{a}\leftarrow\mathbf{0}_{N}$
@@
4:    $\bm{y}\leftarrow\textsc{sequential\_k\_subsets}\left(\beta\left(\bm{X}\bm{q}-\lambda\bm{a}\right),k=2,t\right)$$\triangleright$ Sequential $2$-subsets
@@
9:    $\bm{a}\leftarrow\tau(\bm{y}-\omega\bm{p})+(1-\tau)\bm{a}$triangleright Exponentially weighted average
@@
[…]
@@ 7.1 Metastable state distributions in MNIST
[…]
@@
We start by investigating how often our Hopfield networks converge to metastable states, an important aspect for understanding the network’s dynamics. To elucidate this, we examine $\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{q}^{(t)})$ for the MNIST dataset (LeCun et al., 1998), probing the number of nonzeros in these vectors. We set a threshold $>0.01$ for the softmax method (1-entmax). For the sparse transformations we do not need a threshold, since they have exact retrieval.
@@ 7.2 Hopfield dynamics and basins of attraction
Figure 9: Left: contours of the energy function and optimization trajectory of the CCCP iteration ($\beta=1$) for $\hat{\bm{y}}_{\Psi}(\bm{z})=\frac{\bm{z}}{||\bm{z}||}$. The white regions correspond to infinite energies, due to the hard constraints. Right: attraction basins associated with each pattern ($\beta=10$).
@@
Figure 10: Left: contours of the energy function and optimization trajectory of the CCCP iteration ($\beta=1$) for $\hat{\bm{y}}_{\Psi}(\bm{z})=(\bm{z}-\mu_{\bm{z}})/\sqrt{\sigma_{\bm{z}}^{2}+\epsilon}$. Here we show trajectories for 3D data points and contours according with the $\Psi(\bm{q})$ restrictions mentioned in §3, where the contours lie in the plane $z=-x-y$, intersected with the sphere of radius $\sqrt{D-1}$. White regions correspond to infinite energy. Right: attraction basins associated with each pattern ($\beta=10$).
Figures 8, 9 and 10 illustrate the optimization trajectories and basins of attraction across different queries and artificially generated memory pattern configurations for two families of sparse transformations: $\alpha$-entmax and $\gamma$-normmax. We use the post-transformations $\hat{\bm{y}}_{\Psi}(\bm{z})=\bm{z}$ (identity), $\hat{\bm{y}}_{\Psi}(\bm{z})=\frac{\bm{z}}{||\bm{z}||}$ (normalization) and $\hat{\bm{y}}_{\Psi}(\bm{z})=(\bm{z}-\mu_{\bm{z}})/\sqrt{\sigma_{\bm{z}}^{2}+\epsilon}$ (layer normalization), respectively, which were covered in §3.3. We used $\epsilon=10^{-8}$, as is commonly done in layer normalization for numerical stability. We use $\alpha\in\{1,1.5,2\}$ for $\alpha$-entmax and $\gamma\in\{2,5\}$ for $\gamma$-normmax (where we apply the bisection algorithm described in Appendix A).
@@
In Figure 8, as $\alpha$ increases, $\alpha$-entmax converges more often to a single pattern, whereas $\gamma$-normmax tends to converge towards an attractor which is a uniform average of some patterns. This behavior is also observable in the basins of attraction (right plot), where larger values of $\alpha$ result in fewer regions converging to metastable states. In Figure 9, it is observed that the converged patterns consistently align along the circle with infinite energy outside. This observation is in line with expectations, considering the function $\hat{\bm{y}}_{\Psi}$, derived from $\Psi(\bm{q})=I_{\|.\|\leq r}(\bm{q})$, with $r=1$, that reflects the projected space constraint performed by normalization. In this figure, the local minima of the energy function tend to cluster around a set of memories, whereas in the basins of attraction, the trends closely resemble those in Figure 8, with the exception of the softmax and 1.5-entmax cases, where more attraction areas are present. In Figure 10, we generate synthetic 3D data and plot the contours and trajectories obtained through the Hopfield update rules (CCCP iterations). In this case, we have $\Psi(\bm{q})=I_{S}(\bm{q})$, where $S=\{\bm{q}\mid\|\bm{q}\|\leq\sqrt{D-1}\text{ and }\bm{1}^{\top}\bm{q}=0\}$, which corresponds to layer normalization, and therefore we can represent a query $\bm{q}=(q_{1},q_{2},q_{3})$ in the 2D plane through coordinates $(q_{1},q_{2})$, with $q_{3}=-q_{1}-q_{2}$. After the first iteration, the points converge to the conditions specified by this indicator function, i.e., $\bm{1}^{\top}\bm{q}=0$ plane and $\sqrt{D-1}$ radius sphere. [^6] The $D-1$ term arises because the layer normalization operation uses an unbiased standard deviation. This is a special case of the scenario discussed in §3 where no trainable parameters are present. Similarly to normalization, trajectories tend to converge to clusters of memories, and the basins of attraction exhibit a greater number of attraction areas.
@@ 7.3 Retrieval capacity
We assess the retrieval capacity on three image datasets: MNIST (LeCun et al., 1998), CIFAR10 (Krizhevsky, 2009), and Tiny ImageNet (Le et al., 2015). Prior to inputting the images as queries to the network, we normalize all pixel values to the interval $[-1,1]$. The images are flattened into a single vector when fed to the Hopfield network. During the masking process, pixels outside the mask were set to 0. When introducing Gaussian noise to the images, we ensured that pixel values were clipped, preserving all values within the $[-1,1]$ interval. A query is successfully retrieved when its cosine similarity falls above a predefined threshold of $\epsilon>0.9$. Plotted are the medians of the ratio of successfully retrieved patterns (success retrieval rate) and the interquartile range for 5 runs with different memories for the methods described in Table 2.
In Figure 11 and Figure 12, we can observe that the classic Hopfield networks of Hopfield (1982) and the dense associative models of Krotov and Hopfield (2016) and Demircigil et al. (2017) struggle to successfully retrieve patterns, most noticeable for the former, even for a low number of memories or for low levels of noise. Notably, in modern Hopfield networks with $\beta=1$ (first row of Figure 11), all variants—$\alpha$-entmax and $\gamma$-normmax—show ideal behavior on the MNIST dataset. They also demonstrate accurate behavior on the other datasets, with or without the normalization and layer normalization post-transformations. Additionally, these methods exhibit graceful degradation as the number of stored memories increases. A similar behavior is observed for $\beta=0.1$ (second row), with performance improving as $\alpha$ increases, for $\alpha$-entmax methods, as $\gamma$ decreases for $\gamma$-normmax. One can also see that $2$-entmax with $\hat{\bm{y}}_{\Psi}(\bm{z})={\bm{z}}/{\|\bm{z}\|}$ (normalization) demonstrates even better performance across all datasets, indicating the positive contribution of this specific $\hat{\bm{y}}_{\Psi}(\bm{z})$. Superior performance is also observed with $\hat{\bm{y}}_{\Psi}(\bm{z})=(\bm{z}-\mu_{\bm{z}})/{\sqrt{\sigma_{\bm{z}}^{2}+\epsilon}}$ (layer normalization), although not as good as the former. Similar behavior can be observed in Figure 12 but now in terms of the noise standard deviation. Detailed plots of the HFY networks, using $\hat{\bm{y}}_{\Omega}$ as either $\alpha$-entmax or $\gamma$-normmax for different $\alpha$ and $\gamma$ values and different $\hat{\bm{y}}_{\Psi}$, can be found in Appendix C.1.
@@ 7.4 Sparse and structured transformations in multiple instance learning
Table 4 shows the results. We observe that for MNIST with $K=1$, $1$-entmax outperforms the remaining methods. Normmax shows consistent results across datasets achieving near-optimal performance, arguably due to its ability to adapt to near-uniform metastable states of varying size. We also observe that, for $K>1$, the $k$-subsets approach achieves top performance when $k=K$, as expected. We conjecture that this is due to the ability of SparseMAP with $k$-subsets for $k=K$ to retrieve exactly the $K$ positive instances in the bag, whereas other values of $k$ might under or over-retrieve. In the MIL benchmarks, SparseMAP pooling surpasses sparse pooling variants for 2 out of 3 datasets.
@@
Methods $K$=1 $K$=2 $K$=3 $K$=5 Fox Tiger Elephant
@@ 7.5 Post-transformations in multiple instance learning
$1$-entmax (softmax) Identity $63.6\pm 1.7$ $86.9\pm 1.0$ $91.3\pm 1.0$
@@
$2$-entmax (sparsemax) Identity $63.7\pm 1.7$ $86.8\pm 0.9$ $91.6\pm 0.5$
@@
$5$-normmax Identity $61.9\pm 1.7$ $86.9\pm 1.0$ $\underline{91.9\pm 0.6}$
@@ 7.6 Structured Rationalizers
Sparse Attention sparsemax .82 (.81/.83) .93 (.93/.93) .89 (.89/.90) .019 (.016/.021) 48 (.41/.55)
fusedmax .81 (.81/.82) .92 (.91/.92) .88 (.87/.89) .018 (.017/.019) 39 (.29/.53)
@@ B.2 Proof of Proposition 3
[…]
@@
[…]
@@ B.3 Proof of Proposition 7
To show the second upper bound, use the fact that $E_{\mathrm{concave}}(\bm{q})\leq 0$, which leads to $E(\bm{q})\leq E_{\mathrm{convex}}(\bm{q})=\frac{1}{2}\|\bm{q}-\bm{\mu}_{\bm{X}}\|^{2}+\frac{1}{2}(M^{2}-\|\bm{\mu}_{\bm{X}}\|^{2})=\frac{1}{2}\|\bm{q}\|^{2}-\bm{q}^{\top}\bm{\mu}_{\bm{X}}+\frac{1}{2}M^{2}$. Note that $\|\bm{q}\|=\|\bm{X}^{\top}\bm{y}\|\leq\sum_{i}y_{i}\|x_{i}\|\leq M$ and that, from the Cauchy-Schwarz inequality, we have $-\bm{q}^{\top}\bm{\mu}_{\bm{X}}\leq\|\bm{\mu}_{\bm{X}}\|\|\bm{q}\|\leq M^{2}$. Therefore, we obtain $E(\bm{q})\leq\frac{1}{2}\|\bm{q}\|^{2}-\bm{q}^{\top}\bm{\mu}_{\bm{X}}+\frac{1}{2}M^{2}\leq\frac{1}{2}M^{2}+M^{2}+\frac{1}{2}M^{2}=2M^{2}$.
@@ B.4 Proof of Proposition 9
A stationary point is a solution of the equation $-\nabla E_{\mathrm{concave}}(\bm{q})=\nabla E_{\mathrm{convex}}(\bm{q})$. Using the expression for gradients (B.1), this is equivalent to $\bm{q}=\bm{X}^{\top}\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{q})$. If $\bm{x}_{i}=\bm{X}^{\top}\bm{e}_{i}$ is not a convex combination of the other memory patterns, $\bm{x}_{i}$ is a stationary point iff $\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{x}_{i})=\bm{e}_{i}$. We now use the margin property of sparse transformations (19), according to which the latter is equivalent to $\beta\bm{x}_{i}^{\top}\bm{x}_{i}-\max_{j\neq i}\beta\bm{x}_{i}^{\top}\bm{x}_{j}\geq m$. Noting that the left hand side equals $\beta\Delta_{i}$ leads to the desired result.
@@
where we invoked the Cauchy-Schwarz inequality in the last step. Since the patterns are normalized (with norm $M$), [^7] In fact, the result still holds if patterns are not normalized but have their norm upper bounded by $M$, i.e., if they lie within a ball of radius $M$ and not necessarily on the sphere. we have from the triangle inequality that $\|\bm{x}_{i}-\bm{x}_{j}\|\leq\|\bm{x}_{i}\|+\|\bm{x}_{j}\|=2M$; using the assumption that $\Delta_{i}\geq\frac{m}{\beta}+2M\epsilon$, we obtain ${\bm{q}^{(0)}}^{\top}(\bm{x}_{i}-\bm{x}_{j})\geq\frac{m}{\beta}$, which from the previous points ensures convergence to $\bm{x}_{i}$ in one iteration.

### v5

@@ 6.2 Sequential recall
Algorithm 3 Sequential recall using SparseMAP with sequential 22-subsets. $t>0$ denotes the transition score and $\omega\geq 1$ is a coefficient which promotes sequential order by boosting the emission score of the last recalled pattern.
[…]
@@
4:  $\bm{y}\leftarrow\textsc{sequential\_k\_subsets}\left(\beta\left(\bm{X}\bm{q}-\lambda\bm{a}\right),k=2,t\right)\triangleright$ Sequential 22-subsets
@@
9:  $\bm{a}\leftarrow\tau(\bm{y}-\omega\bm{p})+(1-\tau)\bm{a}$ $\triangleright$ Exponentially weighted average
@@
The algorithm is presented as Algorithm 3. At each step in the outer loop, the structured Hopfield network with the sequential $k$-subsets model is first queried with $\bm{q}$ (which we would like to be close to some pattern, $\bm{q}\approx\bm{x}_{i}$) and returns a pattern association $\bm{y}$—ideally, this is a two-hot vector indicating the index of the cue pattern and the index of an adjacent pattern, e.g., $\bm{x}_{i+1}$; that is, $\bm{y}\approx\bm{x}_{i}+\bm{x}_{i+1}$. In reality it can be a fractional vector satisfying $\mathbf{1}^{\top}\bm{y}=2$. In the ideal scenario we have $\bm{X}^{\top}\bm{y}\approx\bm{x}_{i}+\bm{x}_{i+1}\approx\bm{q}+\bm{x}_{i+1}$. (Note that in this structured Hopfield update we use a penalty similar to the one used in the penalized free recall algorithm, which we will come back to later.) Then, we subtract the query $\bm{q}$ from $\bm{X}^{\top}\bm{y}$—in the ideal scenario above, this should be close to $\bm{q}+\bm{x}_{i+1}-\bm{q}=\bm{x}_{i+1}$. This becomes the query to the inner Hopfield loop, which we expect to yield the attractor $\bm{x}_{i+1}$—the next pattern to be recalled, and the cue for the next step. The penalties are updated with the difference $\bm{y}-\omega\bm{p}$, where we expect $\bm{y}\approx\bm{e}_{i}+\bm{e}_{i+1}$ and $\bm{p}\approx\bm{e}_{i+1}$. If $\omega=1$, this difference would be close to $\bm{e}_{i}$, the indicator of the $i\textsuperscript{th}$ pattern, which will be penalized in subsequent iterations to avoid being retrieved twice. By choosing $\omega$ slightly greater than one, we also add a small bonus to the $(i+1)\textsuperscript{th}$ pattern, which we would like to be retrieved as part of the pattern association in the next step—this avoids memory jumps.
@@ 7.1 Metastable state distributions in MNIST
Table 3: Distribution of metastable state (in $\%$) in MNIST. The training set is memorized and the test set is used as queries.
@@
We start by investigating how often our Hopfield networks converge to metastable states, an important aspect for understanding the network’s dynamics. To elucidate this, we examine $\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{q}^{(t)})$ for the MNIST dataset (LeCun et al., 1998), probing the number of nonzeros in these vectors. We set a threshold >0.01>0.01 for the softmax method (1-entmax). For the sparse transformations we do not need a threshold, since they have exact retrieval.
@@ 7.2 Hopfield dynamics and basins of attraction
Figure 8: Left: contours of the energy function and optimization trajectory of the CCCP iteration ($\beta=1$) for $\hat{\bm{y}}_{\Psi}(\bm{z})=\bm{z}$. Right: attraction basins associated with each pattern ($\beta=10$; a larger $\beta$ is needed to allow for the $1$-entmax to get $T$-close to a single pattern). White sections converge to a metastable state; for $\alpha=1$ we allow a tolerance of $T=.01$).
@@
Figure 10: Left: contours of the energy function and optimization trajectory of the CCCP iteration ($\beta=1$) for $\hat{\bm{y}}_{\Psi}(\bm{z})=(\bm{z}-\mu_{\bm{z}})/\sqrt{\sigma_{\bm{z}}^{2}+\epsilon}$. Here we show trajectories for 3D data points and contours according with the $\Psi(\bm{q})$ restrictions mentioned in §3, where the contours lie in the plane z=xyz=-x-y, intersected with the sphere of radius $\sqrt{D-1}$. White regions correspond to infinite energy. Right: attraction basins associated with each pattern ($\beta=10$).
[…]
@@
In Figure 8, as $\alpha$ increases, $\alpha$-entmax converges more often to a single pattern, whereas $\gamma$-normmax tends to converge towards an attractor which is a uniform average of some patterns. This behavior is also observable in the basins of attraction (right plot), where larger values of $\alpha$ result in fewer regions converging to metastable states. In Figure 9, it is observed that the converged patterns consistently align along the circle with infinite energy outside. This observation is in line with expectations, considering the function $\hat{\bm{y}}_{\Psi}$, derived from $\Psi(\bm{q})=I_{\|.\|\leq r}(\bm{q})$, with $r=1$, that reflects the projected space constraint performed by normalization. In this figure, the local minima of the energy function tend to cluster around a set of memories, whereas in the basins of attraction, the trends closely resemble those in Figure 8, with the exception of the softmax and 1.5-entmax cases, where more attraction areas are present. In Figure 10, we generate synthetic 3D data and plot the contours and trajectories obtained through the Hopfield update rules (CCCP iterations). In this case, we have $\Psi(\bm{q})=I_{S}(\bm{q})$, where $S=\{\bm{q}\mid\|\bm{q}\|\leq\sqrt{D-1}\text{ and }\bm{1}^{\top}\bm{q}=0\}$, which corresponds to layer normalization, and therefore we can represent a query $\bm{q}=(q_{1},q_{2},q_{3})$ in the 2D plane through coordinates $(q_{1},q_{2})$, with $q_{3}=-q_{1}-q_{2}$. After the first iteration, the points converge to the conditions specified by this indicator function, i.e., $\bm{1}^{\top}\bm{q}=0$ plane and $\sqrt{D-1}$ radius sphere. [^7] The D1D-1 term arises because the layer normalization operation uses an unbiased standard deviation. This is a special case of the scenario discussed in §3 where no trainable parameters are present. Similarly to normalization, trajectories tend to converge to clusters of memories, and the basins of attraction exhibit a greater number of attraction areas.
@@ 7.3 Retrieval capacity
We assess the retrieval capacity on three image datasets: MNIST (LeCun et al., 1998), CIFAR10 (Krizhevsky, 2009), and Tiny ImageNet (Le et al., 2015). Prior to inputting the images as queries to the network, we normalize all pixel values to the interval [1,1][-1,1]. The images are flattened into a single vector when fed to the Hopfield network. During the masking process, pixels outside the mask were set to 0. When introducing Gaussian noise to the images, we ensured that pixel values were clipped, preserving all values within the [1,1][-1,1] interval. A query is successfully retrieved when its cosine similarity falls above a predefined threshold of $\epsilon>0.9$. Plotted are the medians of the ratio of successfully retrieved patterns (success retrieval rate) and the interquartile range for 5 runs with different memories for the methods described in Table 2.
[…]
@@ 7.4 Sparse and structured transformations in multiple instance learning
Table 4 shows the results. We observe that for MNIST with $K=1$, 11-entmax outperforms the remaining methods. Normmax shows consistent results across datasets achieving near-optimal performance, arguably due to its ability to adapt to near-uniform metastable states of varying size. We also observe that, for $K>1$, the $k$-subsets approach achieves top performance when $k=K$, as expected. We conjecture that this is due to the ability of SparseMAP with $k$-subsets for $k=K$ to retrieve exactly the $K$ positive instances in the bag, whereas other values of $k$ might under or over-retrieve. In the MIL benchmarks, SparseMAP pooling surpasses sparse pooling variants for 2 out of 3 datasets.
@@
Methods KK=1 KK=2 KK=3 KK=5 Fox Tiger Elephant
@@ 7.5 Post-transformations in multiple instance learning
11-entmax (softmax) Identity $63.6\pm 1.7$ $86.9\pm 1.0$ $91.3\pm 1.0$
@@
22-entmax (sparsemax) Identity $63.7\pm 1.7$ $86.8\pm 0.9$ $91.6\pm 0.5$
@@
55-normmax Identity $61.9\pm 1.7$ $86.9\pm 1.0$ $\underline{91.9\pm 0.6}$
@@ 7.6 Structured Rationalizers
Sparse Attention sparsemax .82 (.81/.83) .93 (.93/.93) .89 (.89/.90) .019 (.016/.021) .48 (.41/.55)
fusedmax .81 (.81/.82) .92 (.91/.92) .88 (.87/.89) .018 (.017/.019) .39 (.29/.53)
@@ 9 Conclusions
Acknowledgments
This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C64500888200000055 (Center for Responsible AI), by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020, by the Dutch Research Council (NWO) VI.Veni.212.228, and by the Champalimaud Foundation.
@@ B.2 Proof of Proposition 3
Let $\Psi(\bm{q})=I_{S}(\bm{q})$ with $S:=\{\bm{q}\,\,:\,\,\|\bm{q}-\bm{\delta}\|\leq\eta\sqrt{D}\,\,\wedge\,\,\bm{1}^{\top}(\bm{q}-\bm{\delta})=0\}.$ We start by showing that, if $f(\bm{q}):=I_{F}(\bm{q})$ with $F:=\left\{\|\bm{q}\|\leq 1\,\,\wedge\,\,\mathbf{1}^{\top}\bm{q}=0\right\}$, we have $(\nabla f^{*})(\bm{z})=\frac{\bm{z}-\mu_{\bm{z}}\mathbf{1}}{\|\bm{z}-\mu_{\bm{z}}\|}$. By definition, we have $f^{*}(\bm{z})=\max_{\bm{q}}\bm{z}^{\top}\bm{q}=-\min_{\bm{q}}-\bm{z}^{\top}\bm{q}$ subject to $\mathbf{1}^{\top}\bm{q}=0$ and $\|\bm{q}\|\leq 1$. Introducing Lagrange multipliers $\mu$ and $\lambda\geq 0$, we obtain the Lagrangian function $L(\bm{q},\mu,\lambda)=-\bm{z}^{\top}\bm{q}+\mu\mathbf{1}^{\top}\bm{q}+\lambda(\|\bm{q}\|-1)$. We have
@@
Now, consider $g(\bm{q}):=I_{G}(\bm{q})$ with $G:=\{\|\bm{q}\|\leq r\,\,\wedge\,\,\mathbf{1}^{\top}\bm{q}=0\}$ for some $r>0$. We can write $g(\bm{q})=f(\bm{q}/r)$, and using the linear transformation property in Table 1, we have $g^{*}(\bm{z})=f^{*}(r\bm{z})$, and therefore $(\nabla g^{*})(\bm{z})=r(\nabla f^{*})(r\bm{z})=r\frac{\bm{z}-\mu_{\bm{z}}\mathbf{1}}{\|\bm{z}-\mu_{\bm{z}}\|}$. When $r=\eta\sqrt{D}$ this becomes $(\nabla g^{*})(\bm{z})=\eta\frac{\bm{z}-\mu_{\bm{z}}\mathbf{1}}{\sigma_{\bm{z}}}$, where $\sigma_{\bm{z}}:=\sqrt{\frac{1}{D}\sum_{i}(z_{i}-\mu_{\bm{z}})^{2}}$. Finally, observe that we can write $\Psi(\bm{q})=g(\bm{q}-\bm{\delta})$. From the translation property in Table 1, we then have $\Psi^{*}(\bm{z})=\bm{\delta}^{\top}\bm{z}+g^{*}(\bm{z})$. This leads to $(\nabla\Psi^{*})(\bm{z})=\bm{\delta}+(\nabla g^{*})(\bm{z})=\mathrm{LayerNorm}(\bm{z};\eta,\bm{\delta})$.
@@ B.3 Proof of Proposition 7
To show the second upper bound, use the fact that $E_{\mathrm{concave}}(\bm{q})\leq 0$, which leads to $E(\bm{q})\leq E_{\mathrm{convex}}(\bm{q})=\frac{1}{2}\|\bm{q}-\bm{\mu}_{\bm{X}}\|^{2}+\frac{1}{2}(M^{2}-\|\bm{\mu}_{\bm{X}}\|^{2})=\frac{1}{2}\|\bm{q}\|^{2}-\bm{q}^{\top}\bm{\mu}_{\bm{X}}+\frac{1}{2}M^{2}$. Note that $\|\bm{q}\|=\|\bm{X}^{\top}\bm{y}\|\leq\sum_{i}y_{i}\|x_{i}\|\leq M$ and that, from the Cauchy-Schwarz inequality, we have $\bm{q}^{\top}\bm{\mu}_{\bm{X}}\leq\|\bm{\mu}_{\bm{X}}\|\|\bm{q}\|\leq M^{2}$. Therefore, we obtain $E(\bm{q})\leq\frac{1}{2}\|\bm{q}\|^{2}-\bm{q}^{\top}\bm{\mu}_{\bm{X}}+\frac{1}{2}M^{2}\leq\frac{1}{2}M^{2}+M^{2}+\frac{1}{2}M^{2}=2M^{2}$.
@@ B.4 Proof of Proposition 9
A stationary point is a solution of the equation $\nabla E_{\mathrm{concave}}(\bm{q})=\nabla E_{\mathrm{convex}}(\bm{q})$. Using the expression for gradients (B.1), this is equivalent to $\bm{q}=\bm{X}^{\top}\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{q})$. If $\bm{x}_{i}=\bm{X}^{\top}\bm{e}_{i}$ is not a convex combination of the other memory patterns, $\bm{x}_{i}$ is a stationary point iff $\hat{\bm{y}}_{\Omega}(\beta\bm{X}\bm{x}_{i})=\bm{e}_{i}$. We now use the margin property of sparse transformations (19), according to which the latter is equivalent to $\beta\bm{x}_{i}^{\top}\bm{x}_{i}-\max_{j\neq i}\beta\bm{x}_{i}^{\top}\bm{x}_{j}\geq m$. Noting that the left hand side equals $\beta\Delta_{i}$ leads to the desired result.
@@
where we invoked the Cauchy-Schwarz inequality in the last step. Since the patterns are normalized (with norm MM), [^8] In fact, the result still holds if patterns are not normalized but have their norm upper bounded by $M$, i.e., if they lie within a ball of radius $M$ and not necessarily on the sphere. we have from the triangle inequality that $\|\bm{x}_{i}-\bm{x}_{j}\|\leq\|\bm{x}_{i}\|+\|\bm{x}_{j}\|=2M$; using the assumption that $\Delta_{i}\geq\frac{m}{\beta}+2M\epsilon$, we obtain ${\bm{q}^{(0)}}^{\top}(\bm{x}_{i}-\bm{x}_{j})\geq\frac{m}{\beta}$, which from the previous points ensures convergence to $\bm{x}_{i}$ in one iteration.